OpenStudy (anonymous):

what is the mass of air within a 83602.3 cm3 balloon in grams?

OpenStudy (anonymous):

at 75 degrees C

OpenStudy (anonymous):

Find the density of air. Note that density p = m/v (note that the Greek character 'rho' is generally used to denote density, it's not actually a p but looks similar). Rearrange the density formula to find the mass based on the density and volume.

OpenStudy (anonymous):

density of air = 1.014 g/L

OpenStudy (anonymous):

(Where m = mass and v = volume).

OpenStudy (mertsj):

it might help if it were true but 1cm^3=1ml not 1 Liter

OpenStudy (mertsj):

And that applies to water at 4 degrees celsius

OpenStudy (anonymous):

Lol oops my bad.

OpenStudy (anonymous):

so how do i find mass?

OpenStudy (anonymous):

volume x density?

OpenStudy (anonymous):

D= M/v so m =dxv

OpenStudy (anonymous):

p = m/v so m = pv Perhaps slightly unrelated, but attached is a list of unit conversions we used in a chemical engineering unit that some may find useful.

OpenStudy (anonymous):

can i just multiply g/L by cm3 or do i convert somehow?

OpenStudy (mertsj):

So if the volume is constant, why did they give you the temperature?

OpenStudy (anonymous):

I'm assuming you're supposed to know/find the density of "air" at 75 degrees Celsius (and assume that "air" is the same universally).

OpenStudy (anonymous):

yeah i got the density fine. 1.014 g/L. but do i have to convert my 83602.3 cm3 ?

OpenStudy (anonymous):

@mertsj to confuse me?

OpenStudy (anonymous):

Keeping it simple, convert either 1.014 g/L to g/cm^3 OR 83602.3 cm^3 to L, depending on which makes the most sense to you (think about which number is larger and what the conversion factor will be). Then multiply.

OpenStudy (anonymous):

it needs to be in grams. so 83.6023 liters and then multiply?

OpenStudy (mertsj):

83602.3 cm^3=83602.3 ml=83.6023 L